OpenStudy (anonymous):

Let F(x)=X^2-16 and G(x)=10-x. Find (F-G)(0)=

hero (hero):

F(x) - G(x) = (x^2 - 16) - (10 - x) = x^2 - 16 - 10 - x = x^2 - x - 26 = 0^2 - 0 - 26 = -26